Part C 0.175 m MgF2 Calculate the boiling point of the solution in part C, assuming complete dissociation.
Answer #1

Tf = -Kf*m*i

i = Mg+2 and 2 F- = 3 ions

kf = -1.86

dTf 0 -1.86*0.175*3 = -0.9765

TF = -0.9765°C